# On Oscillatory Integral Operators Satisfying the cinematic curvature condition

### v1

@@ 5.3. Two-ends: Broad Estimate
In this subsection we prove (5.7). Let $\alpha:=\frac{n-1}{3n-1}$ and thus
@@ 5.3.1. $L^{q}$-estimates using refined decoupling
$u_{2}^{2}R^{\frac{n-2}{4}}|\mathbb{{\bf T}^{\prime}}|^{1/2}\lesssim\max_{\theta\in\Theta_{R^{-1/2}}}\|f_{k,\tau}\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}$
@@
(5.16) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{q}_{A/2}(X)}^{q}\lessapprox_{\phi}R^{-1/2}\mu^{\frac{2}{n-2}}\big(\#\mathbb{T}\big)^{-\frac{1}{n-2}}\|f\|_{2}^{2}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{\,q-2}\\ +{\rm RapDec}(R)\,\|f\|_{2}^{q}$
@@ 5.3.2. A simple $L^{2}$-estimate
Since each $T\in\mathbb{T}$ intersects $O(l)$ cubes $Q$s, Lemma 3.7 yields
(5.18) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{2}_{A/2}(X)}^{2}\lessapprox_{\phi}lR^{1/2}\|f\|_{L^{2}(\mathbb{R}^{n-1})}^{2}+{\rm RapDec}(R)\,\|f\|_{2}^{2}$
@@ 5.3.3. Interpolation
(5.19) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}\lessapprox_{\phi}R^{-\frac{n-3}{12n-4}}l^{\frac{n-3}{6n-2}}\big(\mu\big(\#\mathbb{T}\big)^{-1/2}l^{1/2}\big)^{\frac{2}{3n-1}}+{\rm RapDec}(R)\,\|f\|_{2}^{2}$
@@
(5.20) $\mu\big(\#\mathbb{T}\big)^{-1/2}l^{1/2}\lesssim R^{\varepsilon^{2}}$
Since $l\lesssim R^{1/2}$, (5.7) follows by combining (5.19) and (5.20).
@@ 6. An $\varepsilon$-removal lemma
Suppose $p_{0}\geq 2$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition. We are given that the estimate
(6.1) $\|T^{\lambda}f\|_{L^{p}(B^{n}_{R})}\lesssim_{\phi,a,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
@@
(6.2) $\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{\phi,a,p}\|f\|_{L^{p}(\mathbb{R}^{n-1})}$
@@
[7, Lemma 12.2] Suppose $p_{0}\geq 2$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition. Assume that (6.1) holds uniformly for $1\leq R\leq\lambda$ whenever $p\geq p_{0}$. To prove (6.2) for all $p>p_{0}$ it suffices to show that for all $\varepsilon>0$ the estimate
(6.3) $\|T^{\lambda}f\|_{L^{p_{0}}(S)}\lesssim_{\varepsilon,\phi,a}R^{\varepsilon}\|f\|_{L^{p_{0}}(B^{n-1})}$
@@
The main result of this subsection is stated below and can be deduced by a minor modification of the argument in [7, Lemma 12.3].
@@
(6.4) $\|T^{\lambda}f\|_{L^{p_{0}}(S)}\lesssim_{\varepsilon,\phi,a}R^{\varepsilon}\|f\|_{L^{p_{0}}(B^{n-1})}$
@@
As shown in [7, Proof of Lemma 12.3], it suffices to show that the inequality
@@
where $\{\psi_{D}\}$ is a smooth partition of unity subordinate to this covering. Once we have such a decomposition, (6.6) follows by an application of the Hausdorff–Young inequality together with Hölder’s inequality.

### v2

@@ 5.3. Two-ends: Broad Estimate
In this subsection we prove (5.8). Let
(5.9) $q=\frac{2n}{n-2},\qquad\alpha:=\frac{n-1}{3n-1}$
and thus
@@ 5.3.1. $L^{q}$-estimates using refined decoupling
Define $s$ by
$\frac{\alpha}{2}+\frac{1-\alpha}{s}=\frac{1}{r}$
A direct computation shows that there exists $\gamma\in[0,1]$ with
(5.18) $\frac{1}{s}-\frac{1}{q}+\gamma\big(\frac{1}{2}-\frac{1}{s}\big)=\frac{1}{2}\big(\frac{1}{2}-\frac{1}{q}\big)$
@@
$u_{2}^{2}R^{\frac{n-2}{2}(1-\gamma)}|\mathbb{{\bf T}^{\prime}}|^{\gamma}\lesssim\max_{\theta\in\Theta_{R^{-1/2}}}\|f_{k,\tau}\|_{L_{\mathrm{avg}}^{2}(3\tilde{\theta})}^{2}+{\rm RapDec}(R)\|f\|_{2}^{2}$
@@
(5.19) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{q}_{A/2}(X)}\lessapprox_{\phi}\\ R^{-\frac{n-2}{2}(1-\gamma)\big(\frac{1}{2}-\frac{1}{s}\big)}\mu^{\frac{1}{2}-\frac{1}{q}}\big(\#\mathbb{T}\big)^{\frac{1}{q}-\frac{1}{s}-\gamma\big(\frac{1}{2}-\frac{1}{s}\big)}\|f\|_{2}^{\frac{2}{s}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{s}}\\ +{\rm RapDec}(R)\,\|f\|_{2}$
@@ 5.3.2. A simple $L^{2}$-estimate
Since each $T\in\mathbb{T}$ intersects $O(l)$ cubes $Q$, Lemma 3.7 yields
(5.21) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{2}_{A/2}(X)}\lessapprox_{\phi}l^{1/2}R^{1/4}\|f\|_{L^{2}(\mathbb{R}^{n-1})}+{\rm RapDec}(R)\,\|f\|_{2}$
@@ 5.3.3. Interpolation
$\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}\lessapprox_{\phi}\\ R^{-\frac{n-2}{2}(1-\gamma)\big(\frac{1}{2}-\frac{1}{s}\big)(1-\alpha)+\frac{\alpha}{4}}\mu^{\big(\frac{1}{2}-\frac{1}{q}\big)(1-\alpha)}\big(\#\mathbb{T}\big)^{[\frac{1}{q}-\frac{1}{s}-\gamma\big(\frac{1}{2}-\frac{1}{s}\big)](1-\alpha)}l^{\frac{\alpha}{2}}\\ \|f\|_{2}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}+{\rm RapDec}(R)\,\|f\|_{2}$
From (5.18),
(5.22) $\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}\lessapprox_{\phi}\\ R^{-\frac{n-2}{4}\big(\frac{1}{2}-\frac{1}{q}\big)(1-\alpha)+\frac{\alpha}{4}}\mu^{\big(\frac{1}{2}-\frac{1}{q}\big)(1-\alpha)}\big(\#\mathbb{T}\big)^{-\frac{1}{2}\big(\frac{1}{2}-\frac{1}{q}\big)(1-\alpha)}l^{\frac{\alpha}{2}}\|f\|_{2}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}\\ +{\rm RapDec}(R)\,\|f\|_{2}$
@@
$\mu\big(\#\mathbb{T}\big)^{-1/2}l^{1/2}\lesssim R^{\varepsilon^{2}}$
Combining this with (5.22) and (5.9) we have
$\|T^{\lambda}f_{k}\|_{\textup{BL}^{p}_{A}(X)}\lessapprox_{\phi}R^{-\frac{n-3}{4(3n-1)}}l^{\frac{n-3}{2(3n-1)}}\|f\|_{2}^{\frac{2}{r}}\max_{\theta\in\Theta_{R^{-1/2}}}\|f\|_{L^{2}_{\mathrm{avg}}(3\tilde{\theta})}^{1-\frac{2}{r}}+{\rm RapDec}(R)\,\|f\|_{2}$
Since $l\lesssim R^{1/2}$, this gives (5.8).
@@ 6. An $\varepsilon$-removal lemma
Suppose $2\leq r\leq p_{0}$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition. We are given that the estimate
(6.1) $\|T^{\lambda}f\|_{L^{p}(B^{n}_{R})}\lesssim_{\phi,a,\varepsilon,p}R^{\varepsilon}\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
@@
(6.2) $\|T^{\lambda}f\|_{L^{p}(\mathbb{R}^{n})}\lesssim_{\phi,a,p}\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
@@
[8, Lemma 12.2] Suppose $p_{0}\geq 2$. Let $T^{\lambda}$ be an oscillatory integral operator satisfying the cinematic curvature condition. Assume that (6.1) holds uniformly for $1\leq R\leq\lambda$ whenever $p\geq p_{0}$. To prove (6.2) for all $p>p_{0}$ it suffices to show that for all $\varepsilon>0$ the estimate
(6.3) $\|T^{\lambda}f\|_{L^{p_{0}}(S)}\lesssim_{\varepsilon,\phi,a}R^{\varepsilon}\|f\|_{L^{r}(B^{n-1})}$
@@
The main result of this subsection is stated below and can be deduced by a minor modification of the argument in [8, Lemma 12.3].
@@
(6.4) $\|T^{\lambda}f\|_{L^{p_{0}}(S)}\lesssim_{\varepsilon,\phi,a}R^{\varepsilon}\|f\|_{L^{r}(B^{n-1})}$
@@
As shown in [8, Proof of Lemma 12.3], it suffices to show that
(6.6) $\big(\sum_{j=1}^{N}\|P_{j}f\|_{L^{r}(\mathbb{R}^{n-1})}^{p}\big)^{\frac{1}{p}}\lesssim\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
This estimate follows by interpolating between the endpoint cases $p=r$ and $p=\infty$. By Young’s inequality,
$\|P_{j}f\|_{L^{r}(\mathbb{R}^{n-1})}\leq\|\widehat{\eta}_{R_{1}}\|_{L^{1}(\mathbb{R}^{n-1})}\|f\|_{L^{r}(\mathbb{R}^{n-1})}\lesssim\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
and thus the estimate (6.6) holds when $p=\infty$. It remains to check that
$\big(\sum_{j=1}^{N}\|P_{j}f\|_{L^{r}(\mathbb{R}^{n-1})}^{r}\big)^{\frac{1}{r}}\lesssim\|f\|_{L^{r}(\mathbb{R}^{n-1})}$
This estimate follows via interpolation between the endpoint cases $r=2$ and $r=\infty$. For $r=\infty$ this estimate follows from Young’s inequality. Now we check that this estimate holds for $r=2$. By duality, it suffices to show
@@
where $\{\psi_{D}\}$ is a smooth partition of unity subordinate to this covering. Once we have such a decomposition, (LABEL:7749) follows by an application of the Hausdorff–Young inequality together with Hölder’s inequality.
@@
Now we conclude that the estimate (1.1) holds for $p$ satisfying (1.2) and $r$ satisfying
$\frac{r}{r-1}<\frac{n-2}{n}p$
It remains to verify the boundary case $\frac{r}{r-1}=\frac{n-2}{n}p$. This follows from from the bilinear interpolation argument in [12, 13].